# IGNN-Solver: A Graph Neural Solver for Implicit Graph Neural Networks

### v1

@@ Abstract
Implicit graph neural networks (IGNNs), which exhibit strong expressive power with a single layer, have recently demonstrated remarkable performance in capturing long-range dependencies (LRD) in underlying graphs while effectively mitigating the over-smoothing problem. However, IGNNs rely on computationally expensive fixed-point iterations, which lead to significant speed and scalability limitations, hindering their application to large-scale graphs. To achieve fast fixed-point solving for IGNNs, we propose a novel graph neural solver, IGNN-Solver, which leverages the generalized Anderson Acceleration method, parameterized by a small GNN, and learns iterative updates as a graph-dependent temporal process. Extensive experiments demonstrate that the IGNN-Solver significantly accelerates inference, achieving a 1.5×1.5$\times$1.5 × to 8×8$\times$8 × speedup without sacrificing accuracy. Moreover, this advantage becomes increasingly pronounced as the graph scale grows, facilitating its large-scale deployment in real-world applications.
1 Introduction
@@
Implicit graph neural networks (IGNNs) [20; 33; 7] have emerged as a significant advancement in graph learning frameworks. Unlike traditional graph neural networks (GNNs) that stack multiple explicit layers, IGNNs utilize a single implicit layer formulated as a fixed-point equation. The solution to this fixed-point equation, known as the equilibrium, is equivalent to the output obtained by iterating an explicit layer infinitely. This allows an implicit layer to access infinite hops of neighbors, providing IGNNs with global receptive fields within just one layer [12]. As a result, IGNNs effectively address the long-standing issue of over-smoothing in conventional explicit GNNs and capture of long-range dependencies in graph-structured data [30; 1; 56].
@@
However, existing IGNNs suffer from slow speeds and have difficulty scaling to large-scale graphs [34; 18; 52; 39]. This is primarily because IGNNs derive features by solving for fixed points, demanding substantial computational resources. For example, even on the Citeseer dataset [26] classification task with a small-scale graph, IGNNs require more than $20$ forward iterative computations to nearly converge to a fixed point  [20]. The computational overhead of solving fixed points is amplified by the task scale, resulting in notably slow inference speeds compared to explicit GNNs. This substantial drawback poses challenges for IGNNs in generalizing to large-scale graphs in practical scenarios.
@@
In response to this challenge, we propose a novel graph neural solver for IGNNs, termed IGNN-Solver. It takes the IGNN layer as input, which includes the graph information matrix and layer parameters, and outputs the solution to the corresponding fixed-point equation. Unlike conventional solvers relying on root-finding algorithms like Broyden’s method [2], which compute output features through iterative forward passes in a potentially large GNN, IGNN-Solver offers a distinct advantage by predicting the next iteration step via a tiny graph network. The proposed IGNN-Solver remarkably accelerates the model’s inference speed without compromising accuracy and with only a slight increase in training overhead. This advantage becomes increasingly pronounced as the scale of the graph grows, making it particularly beneficial for deploying IGNNs in large-scale graph tasks.
@@
Our IGNN-Solver comprises two components. First, we introduce a learnable initializer that estimates an optimal initial point for the optimization process. Second, we propose a generalized version of Anderson Acceleration (AA) [2], employing a tiny graph network to model the iterative updates as a sequence of graph-dependent steps. Compared to the solvers proposed for conventional Implicit Neural Networks (INNs) [4; 5; 6], we introduce novel improvements: learning solver parameters through a GNN-based method. This approach circumvents the potential loss of graph information, thereby improving the model’s performance. Moreover, IGNN-Solver has significantly fewer parameters compared to IGNN, and its training is independent of the IGNN’s inference. Consequently, the training of IGNN-Solver proceeds rapidly without sacrificing generalization.
@@
In our experiments, we apply IGNN-Solver to $9$ real-world datasets from diverse domains and scales, including $4$ large-scale datasets: Amazon-all [51], Reddit [21], ogbn-arxiv [22] and ogbn-products [22]. Our results demonstrate that the IGNN-Solver achieves higher accuracy with reduced inference time, showing a 1.5×1.5$\times$1.5 × to 8×8$\times$8 × speedup, and incurs minimal additional training overhead, constituting only $1\%$ of the IGNN training time.
@@
  • We introduce IGNN-Solver, a method designed to predict the next fixed-point iteration of the IGNN layer via a tiny graph network. This innovative approach mitigates the need for extensive iterations, common in conventional IGNNs, thereby substantially accelerating inference while preserving accuracy and minimizing parameter consumption.
  • Compared to conventional solvers proposed for INNs, we have made a novel improvement in our solver by introducing a tiny GNN to learn the parameters. This not only prevents the potential loss of graph information, but also maintains the simplicity and lightweight characteristics of neural solvers.
  • We validate our approach through extensive experiments on $9$ real-world datasets, including $3$ large-scale. Our results demonstrate that IGNN-Solver incurs minimal computational overhead (approximately 1% of the total) and achieves up to a $8$-fold increase in inference speed without compromising accuracy.
2 Related Work
The typical GNN [26] and its variants [42; 21] have been widely used for graph data modeling in various tasks. Different GNNs have been proposed to utilize attention mechanism [42], neighbors sampling [21], pseudo-coordinates [38] and graph fusion [45]. However, due to issues such as over-smoothing, depth, and bottlenecks, these models typically involve finite aggregation layers [30; 1; 56]. To address it, recent works [20; 33; 7] have developed Implicit Graph Neural Networks, encouraging these models to capture long-range dependencies on graphs. Here, we highlight the contributions of our proposed IGNN-Solver through a detailed comparison with Implicit Graph Neural Networks (IGNNs) and Deep Equilibrium Models (DEQs), both of which are closely related to our approach.
Implicit Graph Neural Networks.
@@ 2 Related Work
Although the aforementioned IGNN works well by alleviating the problem of over-smoothing of features by allowing meaningful fixed points to propagate implicitly, the inherent slow inference speed of implicit networks poses a major obstacle to its scalability. The main reason is that the solver of the fixed-point network is inefficient (e.g., the Picard solver used by IGNN [20] and Anderson acceleration solver used by MIGNN [7].), makes the overhead of the fixed-point solver magnified by the task scales. In comparison, our proposed IGNN-Solver accelerates the iteration process of IGNNs, which addresses the most limiting drawback compared to traditional feedforward models.
@@
Fixed-point Solvers for Deep Equilibrium Models.
@@
However, it is well-known that DEQs, as typical implicit models, suffer from slow training and inference speeds [32; 34], which is highly disadvantageous for large-scale scenarios like GraphGPT [58]. To alleviate this issue, recent efforts have explored certain improved solver methods for DEQs, further optimizing root-finding problems and making these models easier to solve: [23] discusses the amortization of the cost of the iterative solver that would otherwise make implicit models slow. [18] proposes a novel gradient estimate for implicit models, named phantom gradient, which significantly accelerates the backward passes in training implicit models. [6] discusses the superiority of learnable solvers over generic solvers in implicit models using a tiny neural network and significantly enhances the efficiency of such models through custom neural solvers.
@@
3 Preliminaries
Explicit GNNs.
Let $\mathcal{G}=\{{\bm{A}},{\bm{X}}\}$ represents an undirected graph, where ${\bm{A}}\in\mathbb{R}^{n\times n}$ is the adjacency matrix indicating the relationships between the nodes in $\mathcal{V}=\{{\bm{v}}_{1},{\bm{v}}_{2},\ldots,{\bm{v}}_{n}\}$, and $n$ is the number of nodes. The node feature matrix ${\bm{X}}=[{\bm{x}}_{1},{\bm{x}}_{2},\ldots,{\bm{x}}_{n}]\in\mathbb{R}^{n\times d}$ contains the features of the nodes, with $d$ representing the feature dimension. Conventional (explicit) GNNs [26; 42] feature a learnable aggregation process centered on the message-passing operation within the graph [19]. This process iteratively propagates information from each node to its neighboring nodes. The formal general structure for each layer $l$ is defined as follows:
@@ 3 Preliminaries
where ${{\bm{H}}^{[l]}}$ represents the hidden node representation, $f_{\theta}$ denotes the parameters in $l$-th layer. A commonly used GNN is Graph Convolutional Network (GCN) [26], defined as ${\bm{H}}^{[l+1]}=\sigma(\hat{{\bm{A}}}{\bm{H}}^{[l]}{\bm{W}}^{[l]})$, where ${{\bm{W}}^{[l]}}$ denotes the weight matrix of the $l$-th layer, $\sigma(\cdot)$ denotes the activation function, ${\hat{{\bm{A}}}}={\tilde{{\bm{D}}}^{-1/2}}({\bm{A}}+{\bm{I}}){\tilde{{\bm{D}}}^{-1/2}}$ represents the symmetric normalized graph matrix. Here, $\tilde{{\bm{D}}}$ is a diagonal matrix with $\tilde{{\bm{D}}}_{ii}=1+\sum\nolimits_{j}{\bm{A}}_{ij}$. GNNs leverage the above message-passing operation in Eq. 1 to learn useful information. Still, they often involve a limited number of $l$ layers due to over-smoothing [1], making it challenging for GNNs to capture the long-range dependency on graphs.
Implicit GNNs.
@@
Similar to traditional explicit GNNs, implicit GNNs [20; 33; 7; 12] also involve an aggregation process, with the distinction that the depth of layers (iteration step $k$) is infinite. The aggregation process in IGNNs is typically defined as ${\bm{Z}}^{[k+1]}=\sigma(\hat{{\bm{A}}}{\bm{Z}}^{[k]}{\bm{W}}+b_{\Omega}({\bm{X}})),k=1,2,\ldots,\infty$, where $b_{\Omega}$ represents affine transformation parameterized by $\bm{\Omega}$, and the weight matrices ${\bm{W}}$ and $\bm{\Omega}$ are globally shared at each iteration step. The IGNN model $f_{\theta}$ is formally described by
${\bm{Z}}^{\star}=f_{{\theta}}({\bm{Z}}^{\star},\hat{{\bm{A}}},{\bm{X}}),$ (2)
@@
where the representation, given as the “internal state” ${\bm{Z}}^{\star}$, is obtained as the fixed-point solution of the equilibrium equation 2. Consequently, the final representation theoretically encompasses information from all neighbors in the graph. In practice, IGNNs capture long-range dependencies within the graph, offering better performance compared to GNNs with finite iterations [20; 33]. Another notable advantage of this framework is its memory efficiency, as it only needs to retain the current state ${\bm{Z}}$ without requiring additional intermediate representations.

### v2

@@ Abstract
Implicit graph neural networks (IGNNs), which exhibit strong expressive power with a single layer, have recently demonstrated remarkable performance in capturing long-range dependencies (LRD) in underlying graphs while effectively mitigating the over-smoothing problem. However, IGNNs rely on computationally expensive fixed-point iterations, which lead to significant speed and scalability limitations, hindering their application to large-scale graphs. To achieve fast fixed-point solving for IGNNs, we propose a novel graph neural solver, IGNN-Solver, which leverages the generalized Anderson Acceleration method, parameterized by a tiny GNN, and learns iterative updates as a graph-dependent temporal process. To improve effectiveness on large-scale graph tasks, we further integrate sparsification and storage compression methods, specifically tailored for the IGNN-Solver, into its design. Extensive experiments demonstrate that the IGNN-Solver significantly accelerates inference on both small- and large-scale tasks, achieving a 1.5×1.5$\times$1.5 × to 8×8$\times$8 × speedup without sacrificing accuracy. This advantage becomes more pronounced as the graph scale grows, facilitating its large-scale deployment in real-world applications. The code to reproduce our results is available at https://github.com/landrarwolf/IGNN-Solver.
Index Terms:
Implicit graph neural networks, large-scale graphs, fixed-point solver, deep equilibrium models.
I Introduction
@@
Implicit graph neural networks (IGNNs) [1, 2, 3] have emerged as a significant advancement in graph learning frameworks. Unlike traditional graph neural networks (GNNs) that stack multiple explicit layers, IGNNs utilize a single implicit layer formulated as a fixed-point equation. The solution to this fixed-point equation, known as the equilibrium, is equivalent to the output obtained by iterating an explicit layer infinitely. This allows the implicit layer to access infinite hops of neighbors, providing IGNNs with global receptive fields within just one layer [4]. As a result, IGNNs effectively address the long-standing issue of over-smoothing in conventional explicit GNNs and capture of long-range dependencies in graph-structured data [5, 6, 7].
@@
However, existing IGNNs suffer from slow speeds and have difficulty scaling to large-scale graphs [8, 9, 10, 11]. This is primarily because IGNNs derive features by solving for fixed points, demanding substantial computational resources. For example, even on the Citeseer dataset [12] classification task with a small-scale graph, IGNNs require more than $20$ forward iterative computations to nearly converge to a fixed point  [1]. The computational overhead of solving fixed points is amplified by the task scale, resulting in notably slow inference speeds compared to explicit GNNs. This substantial drawback poses challenges for IGNNs in generalizing to large-scale graphs in practical scenarios.
@@
In response to this challenge, we propose a novel graph neural solver for IGNNs, termed IGNN-Solver. It takes the IGNN layer as input, which includes the graph information matrix and layer parameters, and outputs the solution to the corresponding fixed-point equation. Unlike conventional solvers, which compute output features through iterative forward passes in a potentially large GNN (relying on root-finding algorithms like Broyden’s method [13]), the IGNN-Solver predicts the next iteration step using a tiny graph network. Consequently, the proposed IGNN-Solver remarkably accelerates the model’s inference speed without compromising accuracy and with only a slight increase in training overhead. This advantage becomes increasingly pronounced as the scale of the graph grows, making it particularly beneficial for deploying IGNNs in large-scale graph tasks.
@@
Our IGNN-Solver comprises two components. First, we introduce a learnable initializer that estimates an optimal initial point for the optimization process. Second, we propose a generalized version of Anderson Acceleration (AA) [13], employing a tiny graph network to model the iterative updates as a sequence of graph-dependent steps. Compared to the solvers proposed for conventional Implicit Neural Networks (INNs) [14, 15, 16], we introduce novel improvements: learning solver parameters through a GNN-based method. This approach circumvents the potential loss of graph information, thereby improving the model’s performance. Moreover, IGNN-Solver has significantly fewer parameters compared to IGNN, and its training is independent of the IGNN’s inference. Consequently, the training of IGNN-Solver proceeds rapidly without sacrificing generalization.
Moreover, to address the challenges posed by large-scale graphs, we incorporate both graph sparsification and storage compression in the solver design. Specifically, to prevent the network from becoming excessively large and deviating from the initial goal of using a compact network for prediction [17, 18], we utilize graph sparsification for a lighter graph. Additionally, directly mapping data from high to extremely low dimensions is not appropriate [19], thereby we employ storage compression technique to mitigate this issue.
@@
In our experiments, we apply IGNN-Solver to nine real-world datasets from diverse domains and scales, including four large-scale datasets: Amazon-all [20], Reddit [21], ogbn-arxiv [22] and ogbn-products [22]. Our results demonstrate that the IGNN-Solver achieves higher accuracy with reduced inference time, showing a 1.5×1.5$\times$1.5 × to 8×8$\times$8 × speedup, and incurs minimal additional training overhead, constituting only $1\%$ of the IGNN training time.
@@
  • We introduce IGNN-Solver, a method designed to predict the next fixed-point iteration of the IGNN layer. This innovative approach mitigates the need for extensive iterations, common in vanilla IGNNs, thereby substantially accelerating inference while preserving accuracy and minimizing parameter consumption.
  • Compared to conventional solvers proposed for INNs, our solver introduces a novel improvement by incorporating a tiny GNN to learn parameters. This approach not only preserves graph information but also maintains the simplicity and lightweight characteristics of neural solvers.
  • For large-scale graph tasks, we integrate custom-designed sparsification and storage compression methods into the solver’s architecture. These tailored techniques enable efficient processing of large-scale graphs, improving performance and reducing overhead.
  • We validate our approach through extensive experiments on nine real-world datasets, including four large-scale. Our results demonstrate that IGNN-Solver incurs minimal computational overhead (approximately 1% of the total) and achieves up to a $8$-fold increase in inference speed without compromising accuracy.
II Related Works
The typical GNN [12] and its variants [23, 21] have been widely used for graph data modeling in various tasks. Different GNNs have been proposed to utilize attention mechanism [23], neighbors sampling [21], pseudo-coordinates [24] and graph fusion [25]. However, due to issues such as over-smoothing, depth, and bottlenecks, these models typically involve finite aggregation layers [5, 6, 7]. To address it, recent works [1, 2, 3] have developed Implicit Graph Neural Networks, encouraging these models to capture long-range dependencies on graphs. Here, we highlight the contributions of our proposed IGNN-Solver through a detailed comparison with Implicit Graph Neural Networks (IGNNs) and Deep Equilibrium Models (DEQs), both of which are closely related to our approach.
II-A Implicit Graph Neural Networks
@@
Although the aforementioned IGNN works well by alleviating the problem of over-smoothing of features by allowing meaningful fixed points to propagate implicitly, the inherent slow inference speed of implicit networks poses a major obstacle to its scalability. The main reason is that the solver of the fixed-point network is inefficient (e.g., the Picard solver used by IGNN [1] and Anderson acceleration solver used by MIGNN [3]), makes the overhead of the fixed-point solver magnified by the task scales. In comparison, our proposed IGNN-Solver accelerates the iteration process of IGNNs, which addresses the most limiting drawback compared to traditional feedforward models.
@@
II-B Fixed-point Solvers for Deep Equilibrium Models
@@
However, it is well-known that DEQs, as typical implicit models, suffer from slow training and inference speeds [35, 8], which is highly disadvantageous for large-scale scenarios like GraphGPT [36]. To alleviate this issue, recent efforts have explored certain improved solver methods for DEQs, further optimizing root-finding problems and making these models easier to solve. [37] discusses the amortization of the cost of the iterative solver that would otherwise make implicit models slow. [9] proposes a novel gradient approximation method for implicit models, which significantly accelerates the backward passes in training implicit models. [16] discusses the superiority of learnable solvers over generic solvers in implicit models using a tiny neural network and significantly enhances the efficiency of such models through custom neural solvers.
@@
III Preliminaries
III-A Explicit GNNs
Let $\mathcal{G}=\{{\bm{A}},{\bm{X}}\}$ represent an undirected graph, where ${\bm{A}}\in\mathbb{R}^{n\times n}$ is the adjacency matrix indicating the relationships between the nodes in $\mathcal{V}=\{{\bm{v}}_{1},{\bm{v}}_{2},\ldots,{\bm{v}}_{n}\}$, and $n$ is the number of nodes. The node feature matrix ${\bm{X}}=[{\bm{x}}_{1},{\bm{x}}_{2},\ldots,{\bm{x}}_{n}]\in\mathbb{R}^{n\times d}$ contains the features of the nodes, with $d$ representing the feature dimension. Conventional (explicit) GNNs [12, 23] feature a learnable aggregation process centered on the message-passing operation within the graph [38]. This process iteratively propagates information from each node to its neighboring nodes. The formal general structure for each layer $l$ is defined as follows:
@@ III-A Explicit GNNs
where ${{\bm{H}}^{[l]}}$ represents the hidden node representation, $f_{\theta}$ denotes the parameters in $l$-th layer. A commonly used GNN is Graph Convolutional Network (GCN) [12], defined as ${\bm{H}}^{[l+1]}=\sigma(\widehat{{\bm{A}}}{\bm{H}}^{[l]}{\bm{W}}^{[l]})$, where ${{\bm{W}}^{[l]}}$ denotes the weight matrix of the $l$-th layer, $\sigma(\cdot)$ denotes the activation function, and ${\widehat{{\bm{A}}}}={\tilde{{\bm{D}}}^{-1/2}}({\bm{A}}+{\bm{I}}){\tilde{{\bm{D}}}^{-1/2}}$ represents the symmetric normalized graph matrix. $\tilde{{\bm{D}}}$ is a diagonal matrix with $\tilde{{\bm{D}}}_{ii}=1+\sum\nolimits_{j}{\bm{A}}_{ij}$. GNNs leverage the above message-passing operation in  (1) to learn useful information. Still, they often involve a limited number of $l$ layers due to over-smoothing [6], making it challenging for GNNs to capture the long-range dependency (LRD) on graphs.
III-B Implicit GNNs
III-B1 Architecture
@@
Similar to traditional explicit GNNs, IGNNs [1, 2, 3, 4] also involve an aggregation process, with the distinction that the depth of layers (iteration step $k$) is infinite. The aggregation process in IGNNs is typically defined as
${\bm{Z}}^{[k+1]}=\sigma({\bm{W}}{\bm{Z}}^{[k]}\widehat{{\bm{A}}}+b_{\Omega}({\bm{X}})),k=1,2,\ldots,\infty,$ (2)
where $b_{\Omega}$ represents affine transformation parameterized by $\bm{\Omega}$, and the weight matrices ${\bm{W}}$ and $\bm{\Omega}$ are globally shared at each iteration step. The IGNN model $f_{\theta}$ is formally described by
${\bm{Z}}^{\star}=f_{{\theta}}({\bm{Z}}^{\star},\widehat{{\bm{A}}},{\bm{X}})=\sigma({\bm{W}}{\bm{Z}}^{\star}\widehat{{\bm{A}}}+b_{\Omega}({\bm{X}})),$ (3)
@@
where the representation, given as the “internal state” ${\bm{Z}}^{\star}$, is obtained as the fixed-point solution of the equilibrium (2). The final representation theoretically encompasses information from all neighbors in the graph. Therefore, IGNNs capture LRD within the graph, offering better performance compared to GNNs with finite iterations [1, 2]. Another notable advantage of this framework is its memory efficiency, as it only needs to retain the current state ${\bm{Z}}$ without requiring additional intermediate representations.
III-B2 Well-posedness
Notably, to obtain a valid representation for any given input ${\bm{X}}$ from IGNNs, we need to solve the unique internal state ${\bm{Z}}^{\star}$ in (3). However, there may not always exist a well-defined solution ${\bm{Z}}^{\star}$ for some inputs ${\bm{X}}$ [1]. To guarantee the existence and uniqueness of the solution to (3), we introduce the concept of well-posedness for IGNNs $f_{\theta}$ with activation function $\sigma$ [1]. Specifically, the weight matrix ${\bm{W}}\in$ $\mathbb{R}^{m\times m}$ and the adjacency matrix $\widehat{{\bm{A}}}\in\mathbb{R}^{n\times n}$ are said to be well-posed for $\sigma$ if the solution ${\bm{Z}}^{\star}\in\mathbb{R}^{m\times n}$ of  (3) exists and is unique for any $b_{\Omega}({\bm{X}})\in\mathbb{R}^{m\times n}$.
Based on the monotone operator theorem [39, 34], a sufficient condition for the well-posedness for IGNNs is defined as follows. Let the nonlinearity $\sigma$ be a non-expansive activation function (e.g., Sigmoid, Tanh, ReLU, etc.), and let $\widehat{{\bm{W}}}=\widehat{{\bm{A}}}^{\top}\otimes{\bm{W}}$.  [^1] We represent the Kronecker product of matrices ${\bm{A}}$ and ${\bm{B}}$ as ${\bm{A}}\otimes{\bm{B}}$. Then, the IGNN model defined in (3) is well-posed as long as ${\bm{I}}-\widehat{{\bm{W}}}\succeq m{\bm{I}}$ for some $m>0$. 222 We represent ${\bm{A}}\succeq{\bm{B}}$ if ${\bm{A}}-{\bm{B}}$ is semi-positive definite. For the case of non-symmetric matrices $\widehat{{\bm{W}}}$, positive definiteness can be defined as the positive definiteness of the symmetric component, i.e., ${\bm{I}}-\widehat{{\bm{W}}}\succeq m{\bm{I}}$ if and only if ${\bm{I}}-\frac{\widehat{{\bm{W}}}+\widehat{{\bm{W}}}^{T}}{2}\succeq m{\bm{I}}$.
It establishes that if $\widehat{{\bm{W}}}$ is symmetric and ${\bm{I}}-\widehat{{\bm{W}}}\succeq m{\bm{I}}$ for some $m>0$, then the existence and uniqueness of the equilibrium point ${\bm{Z}}^{\star}$ are guaranteed, ensuring the model is well-posed. This follows from the sufficient condition that ${\bm{I}}-\widehat{{\bm{W}}}$ is strongly monotone.
To satisfy the well-posedness of IGNNs, a commonly used method [40, 3], is to orthogonally reparameterize ${\bm{W}}$ in the IGNN layer $f_{\theta}$ using the following Cayley map, which is inspired from the unitary operator [41], as
${\bm{W}}=\kappa({\bm{I}}-{\bm{S}})({\bm{I}}+{\bm{S}})^{-1},$ (4)
where $\kappa\in(0,1)$ and ${\bm{S}}={\bm{C}}-{\bm{C}}^{\top}$ is a skew-symmetric matrix with ${\bm{C}}$ as an arbitrary parameterized matrix. From (4), it follows that ${\bm{W}}$ is positive definite and its real eigenvalues are in $(0,1)$. Moreover, based on the definition of $\widehat{{\bm{A}}}$, we can derive the eigenvalues of $\widehat{{\bm{W}}}$ are in $(0,1)$, which ensures that ${\bm{I}}-\widehat{{\bm{W}}}\succeq m{\bm{I}}$, thereby ensuring the well-posedness of IGNN. We refer interested readers to [3] for detailed proof.
III-C Training of IGNNs